# Reasoning-Enhanced Object-Centric Learning for Videos

### v1

@@ 3.2 Slot-based Time-Space Transformer
Table 1: Segmentation results on the MOVi dataset. The first four rows depict the evaluation results for models trained for 100k steps with a batch size of 32. SAVi++* represents results from SAVi++ paper (Elsayed et al., 2022). $*{}^{*}$ denotes the evaluation results for STATM-SAVi++ model trained for 500k steps with a batch size of 64 (Mean $\pm$ standard error over 3 seeds).
@@ A B C D E A B C D E
SAVi-S 62.8 41.6 22.0 6.8 4.0 $91.1$ $70.2$ 50.4 18.4 10.8
STATM-SAVi-S $67.5$ $42.8$ $34.0$ $17.0$ $9.0$ $91.1$ 70.1 $57.7$ $40.9$ $36.9$
@@
STATM-SAVi++ $83.5$ $52.5$ $49.5$ $50.1$ $27.9$ $96.9$ $78.9$ $77.7$ $85.8$ $85.0$
$*{}^{*}$ $76.1\tiny{\pm 0.9}$ $25.8\tiny{\pm 11.3}$ $45.2\tiny{\pm 0.1}$ $48.3\tiny{\pm 0.5}$ $47.1\tiny{\pm 1.3}$ $98.2\tiny{\pm 0.2}$ $48.3\tiny{\pm 15.7}$ $81.9\tiny{\pm 0.2}$ $86.0\tiny{\pm 0.3}$ $84.1\tiny{\pm 0.9}$
$*{}^{*}$ $\bm{85.6}\tiny{\pm 0.6}$ $\bm{60.4}\tiny{\pm 1.2}$ $\bm{52.4}\tiny{\pm 0.2}$ $\bm{57.0}\tiny{\pm 0.4}$ $\bm{55.4}\tiny{\pm 0.9}$ $\bm{98.3}\tiny{\pm 0.2}$ $\bm{84.9}\tiny{\pm 2.5}$ $\bm{82.2}\tiny{\pm 0.2}$ $\bm{89.1}\tiny{\pm 0.2}$ $\bm{88.6}\tiny{\pm 0.5}$
4 Experiments
The central aims of our experiments include: 1) To validate the efficacy of our model, incorporating STATM as a substitute for the transformer encoding block predictor within the SAVi and SAVi++ frameworks. 2) To investigate the effects of varying memory buffer sizes during both the training and inference stages on the performance of the model. 3) To assess the impact of different spatiotemporal methods integrated within STATM on the model’s effectiveness.
Metrics. We selected the Adjusted Rand Index (ARI) (Rand, 1971; Hubert & Arabie, 1985) and the mean Intersection over Union (mIoU) as evaluation metrics. ARI quantifies the alignment between predicted and ground-truth segmentation masks. For scene decomposition assessment, we commonly employ FG-ARI, which is a permutation-invariant clustering similarity metric. It allows us to compare inferred segmentation masks to ground-truth masks while excluding background pixels. mIoU is a widely used segmentation metric that calculates the mean Intersection over Union values for different classes or objects in a segmentation task. It measures the overlap between the predicted segmentation masks and the ground-truth masks, indicating the quality of object segmentation. In the context of video analysis, mIoU is adapted to evaluate the consistency and accuracy of object segmentation and tracking across frames. It provides insights into how well the model captures the spatial relationships between objects in consecutive frames.
Datasets. To evaluate the performance of our model, we utilized the synthetic Multi-Object Video (MOVi) datasets (Research, 2020; Greff et al., 2022). These datasets are divided into five distinct categories: A, B, C, D, and E. MOVi-A and B depict relatively straightforward scenes, each containing a maximum of 10 objects. MOVi-C, D, and E present more intricate scenarios with complex natural backgrounds. MOVi-C, generated using a stationary camera, presents scenes with up to 10 objects. Transitioning to MOVi-D, the dataset extends the object count to accommodate a maximum of 23 objects. Lastly, MOVi-E introduces an additional layer of complexity by incorporating random linear camera movements. Each video sequence is sampled at a rate of 12 frames per second, resulting in a total of 24 frames per second.
Training Setup. We conducted our experiments in JAX (Bradbury et al., 2018) using the Flax (Heek et al., 2020) neural network library. In all experiments except the ablation study in Section 4.2, we used the STAT encoding block in combination with the CS attention architecture, featuring the T+S spatiotemporal fusion approach. For training the STATM-SAVi and SAVi models, we utilized videos comprising of 6 frames at a resolution of 64$\times$64 pixels. The training process is conducted over 100k iterations. Similarly, the STATM-SAVi++ and SAVi++ models were trained on continuous videos consisting of 6 frames at a higher resolution of 128$\times$128 pixels, with training duration encompassing 100k or 500k iterations. The buffer size was unconstrained during training, and the maximum length of effective information was limited to 6 due to the utilization of a 6-frame training sequence. Bounding boxes were used as the conditioning for all models. For more training setup, please refer to Appendix B.
@@ 4 Experiments
Video Refer to caption   Refer to caption  Refer to caption
G.T. Refer to caption   Refer to caption  Refer to caption
SAVi Refer to caption   Refer to caption  Refer to caption
Ours Refer to caption   Refer to caption  Refer to caption
SAVi++ Refer to caption   Refer to caption  Refer to caption
Ours++ Refer to caption   Refer to caption  Refer to caption
Figure 3: Qualitative results of our model compared to SAVi and SAVi++ on the MOVi dataset. Compared with SAVi and SAVi++, our model is slightly better than the SAVi/SAVi++ mode on the relatively simple datasets. As the complexity of the datasets increases, the advantage of our model becomes more pronounced.
4.1 Improvement with STATM
Results. Quantitative results can be seen in Table 1, and qualitative results in Figure 3. In all cases, our models were trained on 6 consecutive frames, but we present results for full (24-frame) sequences at test time. From Table 1, it is evident that our models perform significantly better than the original model. Comparing SAVi and STATM-SAVi, SAVi performs reasonably well on simple datasets but poorly on complex datasets. As the complexity of the datasets increases, the advantage of STATM-SAVi over SAVi becomes more pronounced. In comparison between SAVi++ and STATM-SAVi++, the enhanced model shows a notable improvement, and except for the MOVi-E (due to insufficient training), it can generally match or surpass the optimal results of SAVi++. Further, when considering the best results of $*{}^{*}$ against the official results SAVi++, our model’s performance is markedly superior to the original model, indicating that the model’s benefits do not diminish with increased training. The limitations of depth information and overfitting mentioned in the SAVi++ (Elsayed et al., 2022) do not appear in STATM-SAVi++. The qualitative results from Figure 3 also demonstrate the superior performance of our models (e.g., the notable difference between SAVi and STATM-SAVi in the first frame of MOVi-A). In addition, we provide more qualitative results and visualizations of attention maps for Slot Attention. Please refer to Appendix D for more results.
@@ 4.1 Improvement with STATM
Video Refer to caption  Refer to caption
G.T. Refer to caption  Refer to caption
SAVi++ Refer to caption  Refer to caption
Ours++ Refer to caption  Refer to caption
(c) Object reappears after being occluded
Video Refer to caption
G.T. Refer to caption
SAVi++ Refer to caption
Ours++ Refer to caption
Figure 4: Qualitative results of our model compared to SAVi++. (a) When a new object appears, the original algorithm cannot recognize it, but our model can correctly identifies it after 1-2 frames. (b) and (c) When an object reappears after being obscured, the original algorithm either assigns it to a different slot (color change) or fails to recognize it. In contrast, our model can correctly identify it.
@@ G.T. Refer to caption
In addition, we also tried to use the STATM module to improve the unsupervised scene segmentation model STEVE. The results can be found in the Appendix D.
Generalization. We utilized test sets featuring scenes exclusively comprising previously unseen objects and backgrounds to assess the improved model’s generalization during testing. The results are depicted in Figure 5 (a). Notably, our model exhibits strong generalization capabilities to unfamiliar objects and backgrounds. For additional details, please refer to Appendix D.
Discussion. Certainly, the STATM module, acting as a predictor, significantly enhances the model performance of SAVi and SAVi++. Compared to the original models, the improved model can achieve good performance with fewer training steps and a smaller batch size. STATM-SAVi++ also addresses the overfitting issue of SAVi++ on simple datasets (especially noticeable in MOVi-B). The enhanced models also exhibit good generalization.
4.2 Ablation Study
In this section, we aim to evaluate the influence of different components of STATM, using STATM-SAVi as a baseline. Given the indispensability of the memory module, we focus on two key aspects: 1) The effect of memory buffer size on the model during both training and inference phases; and 2) The influence of different spatiotemporal attention competition and fusion methods on the model.
Ablation Experiment of Memory Module. We have designed two sets of experiments to evaluate the impact of the memory buffer: 1) In the first set, we allowed an unlimited memory buffer length during training, but restricted it to a fixed length during testing, ensuring it didn’t exceed the training buffer’s length. To facilitate evaluation, we have not only assessed the model trained with 6 frames but also extended the training frames to 12, with the 12-frame results available in Appendix E. 2) In the second set, we fixed the buffer length during training, not exceeding the maximum buffer length, and removed any buffer length restrictions during testing. We show results for full (24-frame) sequences during test time in Figure 5 (b) and (c).
@@ 4.2 Ablation Study
Figure 5: Left: (a) Video decomposition results on out-of-distribution evaluation splits with new objects and new backgrounds. Right: Segmentation results with different memory buffers on the MOVi dataset. (b) During training, all available information is utilized with an unlimited buffer length. The $x$-axis represents the model results during testing with buffer sizes ranging from unlimited to limited (6, 4, and 2 frames). (c) Training the model using sub-sequences of 6 frames. The different bars represent the model results during training with buffer sizes of 6, 4, and 2 frames.
Longer-duration video processing presents a challenge to the prediction and inference abilities of the model. It requires that the model extrapolate the learned physical laws of object motion to previously invisible segments. Therefore, the buffer’s role during the testing becomes crucial for inference, especially for object tracking and segmentation beyond the training frame number (see Figure 5 (b)). The prediction module requires additional information to summarize the physical laws of object motion, enabling it to make accurate predictions.
Limiting the buffer length during the training phase reduces the segmentation and tracking capabilities of the model, but the decline is not overly serious. The model’s tracking and segmentation capabilities over a duration equal to the training frames are less affected by memory (see Figure 5 (c)).For more detailed analysis, please refer to Appendix E.
In summary, increasing the memory buffer size during both training and testing phases benefits the improvement of the model’s perceptual capabilities across all datasets. However, for particularly complex datasets like MOVi-E, the excessive increase in the number of training frames may lead to a decline in the model’s segmentation capabilities. In such cases, it might be worth considering improvements to modules like the encoder or corrector to enhance feature extraction capabilities.
Ablation Experiments of Spatiotemporal Fusion and Computation. We conducted ablation experiments of the spatiotemporal fusion method via the CS structure on the MOVi-A dataset. For the ablation experiments related to the spatiotemporal computation structure, we chose the T+S fusion method. Since the AS structure was primarily designed for complex datasets, the computation method ablation experiments were conducted on the MOVi-E dataset. All models were trained using the first 6 frames of the video. The experimental results can be found in Table 2.
Table 2: Result of spatiotemporal fusion method.
@@ A E A E
STATM (CS, ST) 58.4 - 90.9 -
STATM (CS, TS) 61.2 - 89.7 -
STATM (CS, T+S) 67.5 8.5 91.1 36.8
STATM (AS, T+S) - 3.8 - 12.2
On the MOVi-E dataset, the segmentation capability of the AS structure is not as robust as that of the CS structure, but it’s FG-ARI still outperforms the baseline. This suggests the following. 1) Compared to the transform encoding block, it produces more precise predictions for the STATM encoding block with the AS structure as the predictor, enhancing the object segmentation and tracking abilities of slot-based models like SAVi in complex video scenes. 2) As mentioned earlier, the AS structure is designed to handle scenes where objects are not effectively segmented into corresponding slots. Appendix E indicates that with the assistance of initial frame cues, SAVi exhibits decent scene decomposition in the early frames of test videos from the MOVi-E dataset. However, as time progresses, the lack of dynamic temporal interactions among corresponding slots and the impact of complex backgrounds lead to declining segmentation and tracking performance. Currently, models without prompts have limited relevance to our objectives. Hence, we choose not to conduct extensive experiments to verify the capabilities of the STATM with the AS structure.
4.3 Limitations
We used STATM as a prediction module to enhance the perceptual capabilities of slot-based models. However, we didn’t assess our model using real-world datasets. The foundation of our model’s construction is based on the principle that “prediction and correction mutually reinforce each other”. However, our evaluation of the rationality and effectiveness of STATM is based on the experimental results from the correction step, and we haven’t directly tested its physical learning and reasoning abilities. We hope to get a model that performs well in both perception and prediction, which remains an important focus of our future research. In addition, our model may experience a delay (1-2 frames) in recognizing new objects, and the non-immediate release of slots after an object disappears may lead to an insufficient number of slots (more slots needed than the maximum number of objects in the scene).
5 Conclusion
In the real world, all objects follow the laws of physics. Intuitive physics serves as the bridge and connection through which humans comprehend the world. Our research aims to construct biologically plausible deep learning models to explore whether deep learning models can learn physical concepts like humans, and use these learned physical laws to make inferences and predictions about the future motion of objects.

### v2

@@ 4.1. Perception
Table 1. Enhancement results by STATM on models with hints. The first five rows depict the evaluation results for models trained for 100k steps with a batch size of 32. SAVi-SlotFormer denotes our implemented baseline model. SAVi++ represents results from SAVi++ paper (Elsayed et al., 2022). STATM-SAVi++ denotes the evaluation results for STATM-SAVi++ model trained for 500k steps with a batch size of 64 (Mean $\pm$ standard error over 3 seeds).
@@ A B C D E A B C D E
SAVi 62.8 41.6 22.0 6.8 4.0 $91.1$ 70.2 50.4 18.4 10.8
SAVi-SlotFormer 63.5 - - - 7.5 86.4 - - - 31.2
STATM-SAVi (Ours) $67.5$ $42.8$ $34.0$ $17.0$ $9.0$ $91.1$ $70.7$ $57.7$ $40.9$ $36.9$
@@
STATM-SAVi++ (Ours) $83.5$ $52.5$ $49.5$ $50.1$ $27.9$ $96.9$ $78.9$ $77.7$ $85.8$ $85.0$
SAVi++ $76.1{\pm 0.9}$ $25.8{\pm 11.3}$ $45.2{\pm 0.1}$ $48.3{\pm 0.5}$ $47.1{\pm 1.3}$ $98.2{\pm 0.2}$ $48.3{\pm 15.7}$ $81.9{\pm 0.2}$ $86.0{\pm 0.3}$ $84.1{\pm 0.9}$
STATM-SAVi++ (Ours) $\bm{85.6}{\pm 0.6}$ $\bm{60.4}{\pm 1.2}$ $\bm{52.4}{\pm 0.2}$ $\bm{57.0}{\pm 0.4}$ $\bm{55.4}{\pm 0.9}$ $\bm{98.3}{\pm 0.2}$ $\bm{84.9}{\pm 2.5}$ $\bm{82.2}{\pm 0.2}$ $\bm{89.1}{\pm 0.2}$ $\bm{88.6}{\pm 0.5}$
@@
Video
Refer to caption   Refer to caption  Refer to caption
G.T.
Refer to caption   Refer to caption  Refer to caption
SAVi
Refer to caption   Refer to caption  Refer to caption
Ours
Refer to caption   Refer to caption  Refer to caption
SAVi++
Refer to caption   Refer to caption  Refer to caption
Ours++
Refer to caption   Refer to caption  Refer to caption
Figure 3. Qualitative results of our model compared to SAVi and SAVi++ on the MOVi dataset. Compared with SAVi and SAVi++, our model is slightly better than the SAVi/SAVi++ mode on the relatively simple datasets. As the complexity of the datasets increases, the advantage of our model becomes more pronounced.
Table 2. Enhancement Results (FG-ARI%) on STEVE.
Model MOVi-D MOVi-E
STEVE 47.67 52.15
STATM-STEVE (Ours) $\bm{51.73}$ $\bm{55.78}$
Results. Quantitative results can be seen in Table 1 and  2, and qualitative results in Figure 3. From Table 1 and  2, it is evident that STATM can significantly enhance the performance of existing state-of-the-art models. Comparing SAVi, SAVi-SlotFormer and STATM-SAVi, SAVi performs reasonably well on simple datasets but struggles with complex datasets. When SlotFormer is used as the predictor, it enhances SAVi’s performance on complex datasets, yet it reduces its effectiveness on simpler datasets. Conversely, when STATM serves as the predictor, it not only improves SAVi’s performance on complex datasets but also maintains its performance on simple datasets. In comparison between SAVi++ and STATM-SAVi++, the enhanced model shows a notable improvement, and except for the MOVi-E (due to insufficient training), it can generally match or surpass the optimal results of SAVi++. Further, when considering the best results of STATM-SAVi++ against the official results SAVi++, our model’s performance is markedly superior to the original model, indicating that the model’s benefits do not diminish with increased training. The limitations of depth information and overfitting mentioned in the SAVi++ (Elsayed et al., 2022) do not appear in STATM-SAVi++. The qualitative results from Figure 3 also demonstrate the superior performance of our models (e.g., the notable difference between SAVi and STATM-SAVi in the first frame of MOVi-A).
In addition, we also tried to use the STATM module to improve the unsupervised scene segmentation model STEVE. The results can be found in Table 2. We found that STAM remains effective for unsupervised object-centric learning model.
@@
Video
Refer to caption  Refer to caption
G.T.
Refer to caption  Refer to caption
SAVi++
Refer to caption  Refer to caption
Ours++
Refer to caption  Refer to caption
Figure 4. Qualitative results of our model compared to SAVi++. (a) When a new object appears, the SAVi++ cannot recognize it, but our model can correctly identifies it after 1-2 frames. (b) When an object reappears after being obscured, the SAVi++ either assigns it to a different slot (color change) or fails to recognize it. In contrast, our model can correctly identify it.
@@
Memory Cost and Inference Time. The memory cost and inference time (taken to process 250 videos) are listed in Table 3. We can see that adding STATM to the model doesn’t bring much extra memory cost, inference time and parameters. Detailed comparison of parameters can be found in Appendix Table S1 Appendix Section C.
Table 3. Memory cost and inference time on MOVi-A and E validation sets, with each set containing 250 videos, each video contains 24 frames (batch size of 32, on one A100 GPU).
Model Memory (GB) Infer. Time (s)
A E A E
SAVi 25.91 56.42 78.4 141.3
SAVi-SlotFormer 25.92 56.43 102.6 204.1
STATM-SAVi (Ours) 25.91 56.42 81.4 146.4
Generalization.We selected the models trained with a batch size of 32 and 100k training steps to assess its generalization. The test sets utilized the default test split of MOVi-C, D and E dataset, featuring scenes exclusively consist of held-out objects and background images to evaluate generalization. The results are presented in Table 4 and Figure 6 (a).
Table 4. Generalization results on MOVi datasets.
@@ A E A E
C D E C D E
SAVi-S (IID) 22.0 6.8 4.0 50.4 18.4 10.8
SAVi-S (OOD) 21.1 6.3 3.7 52.8 19.8 9.9
STATM-SAVi-S (IID) 34.0 17.0 9.0 57.7 40.9 36.9
STATM-SAVi-S (OOD) 33.2 17.0 8.2 59.7 43.4 35.9
SAVi++ (IID) 47.8 43.6 26.1 76.3 81.5 81.7
SAVi++ (OOD) 46.9 44.0 25.5 77.7 82.2 82.5
STATM-SAVi++ (IID) 49.5 50.1 27.9 77.7 85.8 85.0
STATM-SAVi++ (OOD) 48.7 50.3 27.4 78.9 86.4 85.8
Discussion. Certainly, the STATM module, acting as a predictor, significantly enhances the model performance of SAVi and SAVi++. Compared to the original models, the improved model can achieve good performance with fewer training steps and a smaller batch size. STATM-SAVi++ also addresses the overfitting issue of SAVi++ on simple datasets (especially noticeable in MOVi-B). The enhanced models also exhibit good generalization. Importantly, the integration of a STAT encoding block does not lead to a significant increase in memory cost and inference time.
4.2. Prediction and VQA
In this section, our primary objective is to succinctly validate the performance of STATM in downstream prediction and VQA tasks.
Perception on CLEVRER. Due to the effectiveness of object-centric predictive models largely depends on the quality of slots extracted by upstream perceptual modules. Additionally, MOVi validation sets only includes sequences of 24 frames in length. Therefore, in Table 5, we demonstrate a comparison of perceptual effects on CLEVER with longer sequences of 128 frames. It is evident that STATM-SAVi achieves a significantly higher FG-ARI compared to SAVi, indicating that STATM provides a more pronounced enhancement on SAVi in longer sequences on relatively simple datasets.
Table 5. Perception results on CLEVRER.
[…]
SAVi 0.24 91.4 77.6
STATM-SAVi (Ours) 0.15 93.5 77.6
Prediction. Table 6 displays the evaluation results of visual quality and object dynamics on CLEVRER. It is evident that when using slots extracted by SAVi to train our STATM for prediction, the model shows improvement in all metrics except PSNR, indicating a certain advantage of our model in prediction tasks. When using slots extracted by STATM-SAVi to train STATM for predictions, both video quality and predictive metrics see further enhancements, particularly FG-ARI. This further demonstrates the advantages of our model in both perception and prediction.
Table 6. Evaluations of visual quality (columns 2-4) and object dynamics (columns 5-6) on CLEVRER. SA+ST and ST+ST respectively represent the results of using SAVi and STATM-SAVi for slot extraction followed by training STATM.
[…]
SAVi-Dyn 29.77 0.89 0.19 64.32 18.25
SlotFormer 30.21 0.89 0.11 63.00 49.40
SA+ST(Ours) 30.10 0.89 0.11 63.11 49.55
ST+ST(Ours) 30.22 0.89 0.11 64.56 49.57
Figure 5. Results of long-sequence prediction on CLEVRER. After surpassing a certain time point, results generated by SlotFormer in prediction clearly begin to deviate from the ground truth, exhibiting artifacts such as blurry (orange boxes), incorrect dynamics (red boxes), and inaccurate colors (green boxes). Meanwhile, our model demonstrates good performance.
Figure 5 displays generation results of long-sequence prediction on CLEVRER. We observe that SlotFormer performs well in shorter sequence predictions. However, as time progresses, SlotFormer’s performance deteriorates significantly, exhibiting blurriness, incorrect dynamics, and inaccurate colors, even becoming completely inconsistent with the ground truth. In contrast, our model continues to perform well.
VQA. In Table 7, we present the accuracy on predictive questions. Notably, also as an unsupervised predictive model, our method surpasses the previous state-of-the-art SlotFormer (Wu et al., 2023). Furthermore, on the publicly available CLEVRER leaderboard for the predictive question subset, our approach achieved first rank in the per option setting and second rank in the per question setting.
Table 7. Predictive VQA accuracy on CLEVRER.
Model per opt.(%) per ques.(%)
DCL 90.52 82.03
VRDP 95.68 91.35
SlotFormer 96.50 93.29
STATM (Ours) 96.62 93.63
Discussion. Clearly, our model demonstrates distinct advantages in perception over longer sequences, as well as in downstream prediction and VQA tasks.
4.3. Ablation Study
[…]
@@ 4.3. Ablation Study
Figure 6. (a) Results on out-of-distribution evaluation splits with new objects and backgrounds. (b) Ablation study of the buffer size on model performance during testing phase. The $x$-axis represents buffer sizes during testing. (c) Ablation study of the impact of buffer size on model performance during training phase. The different bars represent buffer sizes.
Table 8. Ablation study of buffer on CLEVRER during the testing phase. ‘Perception’ represents perceptual results of STATM-SAVi with different buffer sizes. ‘Prediction’ shows prediction results of STATM trained with slots extracted by STATM-SAVi with different buffer sizes. ‘VQA’ indicates our model’s accuracy on predictive questions when using various buffer sizes.
Model Perception Prediction VQA
FG-ARI(%) FG-mIoU(%) FG-ARI(%) FG-mIoU(%) per opt.(%) per ques.(%)
STATM (24) 93.5 77.6 64.6 49.6 96.62 93.63
STATM (32) 93.5 77.6 64.5 49.5 96.41 93.09
STATM (48) 93.5 77.6 64.6 49.6 96.07 92.67
STATM (128) 93.5 77.6 64.2 49.3 96.21 92.90
In this section, we aim to evaluate the influence of different components of STATM, using STATM-SAVi as a baseline.
Memory Buffer. We have designed two sets of experiments to evaluate the impact of the memory buffer: 1) In the first set, we allowed an unlimited memory buffer length during training, but restricted it to a fixed length during testing. 2) In the second set, we fixed the buffer length during training, and removed any buffer length restrictions during testing. To facilitate evaluation, we have not only assessed the model trained with 6 frames but also extended the training frames to 12. We show the results in Figure 6 (b) and (c).
Training phase. From Figure 6 (c), we observe that during the perception training phase, changes in buffer size have minimal impact on the model’s performance on simple datasets. For complex datasets, initially increasing the buffer size improves model performance, but further increases eventually lead to a decline. Thus, during the perception training phase, buffer size can be set to 6.
Testing phase. Since the MOVi validation set contains only 24 frames, it does not adequately demonstrate the impact of memory size during the testing phase. Therefore, in Table 8 perception column, we present evaluation results on CLEVRER dataset with a sequence of 128 frames during testing phase. We observe that increasing the buffer size during the testing phase initially improves perceptual outcomes, but beyond a certain size, it has negligible impact. Thus, we set the buffer size to 24 during the testing phase to avoid unnecessary computational costs.
Table 8 prediction and VQA column display the effects of using different buffer sizes of STATM-SAVi to extract slots during the testing phase on prediction and VQA performance. It is observed that slots extracted by SAVi-STATM using different buffer sizes have almost no impact on STATM’s predictive capabilities. The impact of buffer size on VQA is minor, but beyond a certain threshold, VQA performance actually deteriorates. Therefore, for prediction and VQA tasks, we use 24 buffer size of STATM-SAVi to extract slot.
In summary, the buffer size can be set to 6 during the training of both the perception model and the long-term reasoning component. When testing the perception model, the buffer size should be set to 24. For the long-term reasoning component during use, a buffer size of 6 yields good results.
Table 9. Ablation study of perception for different spatiotemporal attention computation and fusion methods.
@@ A E A E
STATM-SAVi (CS, ST) 58.4 - 90.9 -
STATM-SAVi (CS, TS) 61.2 - 89.7 -
STATM-SAVi (CS, T+S) 67.5 8.5 91.1 36.8
STATM-SAVi (AS, T+S) - 3.8 - 12.2
Spatiotemporal Fusion and Computation. In Table 9, we display results for STATM-SAVi using different spatiotemporal attention computation and fusion methods. From the first three rows of the table, it can be seen that the T+S spatiotemporal fusion method is relatively superior to both ST and TS. The last two rows indicate that the CS spatiotemporal attention structure is superior to AS. This may be due to the AS structure causing slot confusion from excessive slot interactions. Therefore, we use STAT with CS and T+S for our experiments by default.
4.4. Limitations
We did not assess our model using real-world datasets. Our perception and prediction models are trained separately, although they share a common predictive structure. In the future, we are more interested in implementing integrated training of our models.
5. Conclusion